# MonoGSDF: Exploring Monocular Geometric Cues for Gaussian Splatting-Guided Implicit Surface Reconstruction

### v3

@@ 3.3 Optimization
Training our Gaussian-guided SDF presents unique challenges: enforcing all Gaussians to satisfy the condition $\bar{s}_{i}=0$ degrades rendering quality by removing opacity variations, while sparse Gaussian-based SDF supervision lacks spatial coverage. Volume rendering methods like 3DGSR [24] introduce computational overhead and limited 2D supervision is insufficient for accurate 3D representation. Our solution introduces a novel framework that enhance the Gaussians with an off-the-shelf geometry estimator and supervise the SDF with a dense 3D regularization through depth back-projection.
Geometry Cues. To enhance reconstruction quality in real-world scenarios, we integrate an off-the-shelf monocular depth estimator [^1] We use DepthAnything V2 [41] in this work. to obtain dense per-pixel disparity priors as $\hat{Z}$. However, these estimated pseudo disparity measurements exhibit significant scale ambiguity and suffer from multi-view inconsistency. While DN Splatter [34] addresses this issue through direct depth alignment, applying a transformation $\bar{D}=s\hat{D}+t$, their method fails to account for shifts in disparity $\hat{Z}$, resulting in distance-dependent scale variations. To ensure robust relative scale consistency, we first utilize the rendered depth to correct the pseudo disparity and obtain the pseudo depth $\hat{D}=a/(s*\hat{Z}+t)+b$:
$\mathcal{L}_{D}=\sum|({a/(s*\hat{Z}+t)+b-\bar{D}}|.$ (7)
where $s,t$ and $a,b$ represent learnable scale and shift parameters. We further convert the aligned depth into pseudo normal $D^\mathbf{\hat{N}}$$\nabla\hat$$|\nabla\hat{D}$$|$, providing a more stable supervision signal for Gaussian training:
$\mathcal{L}_{N}=\sum(1-\mathbf{\hat{N}}\cdot\mathbf{\bar{N}}).$ (8)
where $\mathbf{\bar{N}}$ is rendered Gaussian normal. Therefore, the overall geometry regularization is $\mathcal{L}_{geo}=\lambda_{D}\mathcal{L}_{D}+\lambda_{N}\mathcal{L}_{N}$.
@@
Multi-View Consistent 3D Regularization. Directly supervising SDF with monocular geometry cues like MonoSDF [46] suffers from scale ambiguity and lacks multi-view consistency. Further, volume rendering-based supervision as proposed in 3DGSR [24] and GSDF [45] are inefficient and only supervise SDF with rendered Gaussian depth map in 2D. Our approach combines monocular cues with the inherent advantages of 3D Gaussians that the Gaussian-rendered depth $\bar{D}$ naturally maintains multi-view consistency and enables reliable 3D signal supervision for SDF through back-projection.
More specifically, we first sample $M$ pixels $\{\mathbf{u}=\mathbf{u}_{i}\ |\ \mathbf{u}_{i}\in\mathbb{R}^{2},{\text{for}\ i=1,2,...,M\}}$ from the rendered Gaussian depth map $\bar{D}$. For each ray casts from a pixel, we sample $K=K_{n}+K_{f}$ points $\{t_{k}\}$ from the camera center to the surface, where $K_{n}$ means near surface samples and $K_{f}$ means free space samples. SDF values of each point can be queried as $\{\bar{s}_{k}\}$. For near surface samples within the truncation region $tr$, i.e. $S_{tr}=\{|\bar{D}(\mathbf{u})-t_{k}|\leq tr\}$, we use the distance between the sampled point $t_{k}$ and its surface $\bar{D}(\mathbf{u}_{j})$ as an approximation of ground truth SDF value for supervision:
[…]
For points that are far from the surface $S_{fs}=\{\bar{D}(\mathbf{u})-t_{k}>tr\}$, we apply a free-space supervision as:
@@
Multi-Resolution Regularization. Unlike vanilla 3DGS [16] that uses L1 and SSIM losses with opacity-based pruning, our method links opacity to spatial position, lifting the need for Opacity Resets but potentially causing Gaussian redundancy. While low-to-high frequency strategies [52] can accelerate convergence, they risk 3D structural blurring. Instead, we propose multi-resolution supervision through wavelet transforms, progressively increasing resolution while preserving structural integrity:
${\mathbf{C}_{l}}=\mathcal{W}({\mathbf{C}},l)$ (11)
where $\mathbf{C}$ is ground truth color images, $\mathcal{W}$ represents wavelet transform and $l$ denotes the level of transformation. This approach avoids the zigzag artifacts of direct image rescaling, ensuring stable detail refinement during training. It enables the Gaussian model to first capture a rough scene representation with fewer Gaussians and later refine scene details for a more precise Gaussian distribution. Therefore, the photometric regularization is $\mathcal{L}_{p}=0.8*|{\mathbf{C}_{l}}-\bar{\mathbf{C}}|+0.2*SSIM({\mathbf{C}_{l}},\bar{\mathbf{C}})$.
Objective Function. During training, we jointly optimize explicit Gaussians and implicit SDF. We apply distortion $\mathcal{L}_{d}$ and depth-normal $\mathcal{L}_{n}$ regularization as defined in GOF [48]. The overall objective function is defined as:
$\mathcal{L}=\mathcal{L}_{p}+\mathcal{L}_{geo}+\mathcal{L}_{sdf}+\lambda_{d}\mathcal{L}_{d}+\lambda_{n}\mathcal{L}_{n}.$ (12)
Pruning. In our framework, where opacity is intrinsically linked to spatial position through the SDF relationship, we implement a geometry-aware pruning strategy that removes Gaussians exceeding the truncation threshold: $\bar{s}_{i}>tr$. This approach replaces conventional opacity-based pruning with a physically-grounded criterion that better aligns with our surface-aware representation.
Mesh Extraction. Previous Marching Cubes-based approaches [45, 24] must densely search both occupied and free space, leading to inefficiencies and resolution limitations, especially in unbounded scenes with coordinate distortions. In contrast, our method adapts Marching Tetrahedral [18] and exploits 3DGS’s spatial cues to constrain the search to regions only near Gaussians, avoiding unnecessary exploration of empty space while achieving superior surface alignment and reconstruction efficiency.
@@ 4.1 Experimental Settings
Datasets. We comprehensively evaluate our method on three public datasets: DTU dataset [13] which consists of indoor object scans; Tanks and Temples [17] which features six unbounded outdoor scenes; and Mip-NeRF 360 [3] for measuring novel view synthesis quality.
Baselines. We compare our proposal against several state-of-the-art methods. Among implicit methods, we compare against NeRF [25], VolSDF [44], NeuS [37], N-angelo [21], GeoNeus [10], Instant NGP [27] and MipNeRF 360 [3]. As for explicit Gaussian methods, we compare against 3DGS [16], SuGaR [11], Mip-Splatting [47], 2DGS [12], GSurfel [8], 3DGSR [24], GS-Pull [55] and GOF [48].
Metrics. To evaluate our method, we follow common practice and report surface accuracy as Chamfer Distance and F1-score on DTU and Tanks and Temples, respectively. We measure the visual fidelity of the synthesized novel views with PSNR, SSIM and LPIPS [54] on Mip-NeRF 360.
Implementation. We perform single GPU training (NVIDIA 3090) and use by default $\lambda_{d}=100,\lambda_{n}=0.05,\lambda_{ns}=1000,\lambda_{fs}=10,\lambda_{depth}=0.05,\lambda_{normal}=0.1$ and $\beta=100$. We train 30000 iterations like other methods and do not require extra training for SDF. For more implementation details, please refer to supp. mat..
24 37 40 55 63 65 69 83 97 105 106 110 114 118 122 Mean
NeRF [25] 1.90 1.60 1.85 0.58 2.28 1.27 1.47 1.67 2.05 1.07 0.88 2.53 1.06 1.15 0.96 1.49
VolSDF [44] 1.14 1.26 0.81 0.49 1.25 0.70 0.72 1.29 1.18 0.70 0.66 1.08 0.42 0.61 0.55 0.86
NeuS [37] 1.00 1.37 0.93 0.43 1.10 0.65 0.57 1.48 1.09 0.83 0.52 1.20 0.35 0.49 0.54 0.84
N-angelo [21] 0.49 1.05 0.95 0.38 1.22 1.10 2.16 1.68 1.78 0.93 0.44 1.46 0.41 1.13 0.97 1.07
SuGaR [11] 1.47 1.33 1.13 0.61 2.25 1.71 1.15 1.63 1.62 1.07 0.79 2.45 0.98 0.88 0.79 1.33
2DGS [12] 0.48 0.91 0.39 0.39 1.01 0.83 0.81 1.36 1.27 0.76 0.70 1.40 0.40 0.76 0.52 0.80
GSurfel [8] 0.66 0.93 0.52 0.41 1.06 1.14 0.85 1.29 1.53 0.79 0.82 1.58 0.45 0.66 0.53 0.88
3DGSR [24] 0.68 0.84 0.70 0.39 1.16 0.87 0.77 1.48 1.26 0.87 0.69 0.80 0.42 0.64 0.60 0.81
GS-Pull [55] 0.51 0.56 0.46 0.39 0.82 0.67 0.85 1.37 1.25 0.73 0.54 1.39 0.35 0.88 0.42 0.75
GOF [48] 0.50 0.82 0.37 0.37 1.12 0.74 0.73 1.18 1.29 0.68 0.77 0.90 0.42 0.66 0.49 0.74
Table 1: Quantitative Comparison on the DTU Dataset [13]. We show the Chamfer distance. Our method achieves the highest reconstruction accuracy among all methods. For N-angelo [21], we report the results from UniSDF [35] reproduction and we show the vanilla results in supp. mat..
Figure 3: Surface Reconstruction on the DTU Dataset [1]. We show normal maps rendered from the reconstructed meshes. 2DGS [12] produces overly smooth surfaces with floating artifacts and outliers, while GOF [48] struggles to reconstruct flat surface. In contrast, our method generates smooth yet detailed surfaces without floating artifacts, demonstrating its superior ability to represent geometry accurately.
4.2 Geometry Evaluation
We evaluate our method on the DTU dataset [13], comparing it with state-of-the-art surface reconstruction techniques. As shown in Table 1, our approach achieves the lowest Chamfer Distance, surpassing all existing 3DGS-based and implicit methods in both quality and efficiency. For N-angelo [21], we report results reproduced in [35], as the original values could not be verified. Figure 3 presents qualitative comparisons, where normal maps reveal that 2DGS [12] and GOF [48] often produce noisy, incomplete surfaces with artifacts, particularly struggling with flat regions. Our method mitigates these issues by constraining Gaussians to the surface via SDF, yielding smoother and more accurate reconstructions, especially in challenging areas with sparse camera coverage.
@@ 4.2 Geometry Evaluation
NeuS [37] GeoNeus [10] N-angelo [21] SuGaR [11] 2DGS [12] GOF [48] Ours
@@
Average 0.38 0.35 0.50 0.19 0.30 0.46 0.47
Table 2: Quantitative Results on the Tanks and Temples Dataset [17]. Reconstructions are evaluated with the official evaluation scripts and we report F1-score. Our method outperforms all 3DGS-based surface reconstruction methods and performs comparably with the SOTA neural implicit methods while optimizing significantly faster.
We then compare our method on the Tanks and Temples dataset [17]. The evaluation is focused exclusively on foreground objects, as the ground truth point cloud does not cover background regions. As shown in Table 2, our method achieves results on par with leading implicit approaches [21] while significantly reducing training time—optimizing large unbounded scenes in around 3 hours versus over 12 hours for implicit methods. Additionally, we slightly outperform GOF [48] while maintaining comparable rendering efficiency and faster mesh extraction (Table 4). Figure 4 showcases qualitative results, where our method excels at reconstructing flat surfaces, such as the Barn rooftop, and handling reflective and transparent windows in the Caterpillar and Truck scenes—areas where 2DGS [12] and GOF [48] struggle due to limited views and scene complexity. Our Gaussian-guided implicit surface reconstruction method ensures smooth and accurate reconstructions in these challenging scenarios.
Figure 4: Surface Reconstruction on the Tanks and Temples Dataset [17]. We show the rendered normal maps from reconstructed meshes. Resuls of 2DGS [12] and GOF [48] struggles to reconstruct smooth surfaces around transparent windows and produces noisy results in regions with sparse views. In contrast, our approach effectively handles transparent areas and achieves smooth reconstruction even in sparsely viewed regions.
Figure 5: Surface Reconstruction on the Mip-NeRF 360 Dataset [3]. We show the rendered normal maps from reconstructed meshes. GOF’s [48] meshes are noisy and fail to reconstruct smooth surface where there are reflection and curvature surface. In contrast, our method can generate smoother and detailed surface without any outlier points.
Outdoor Scene Indoor Scene
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
NeRF [25] 21.46 0.458 0.515 26.84 0.790 0.370
Instant NGP [27] 22.90 0.566 0.371 29.15 0.880 0.216
MipNeRF360 [3] 24.47 0.691 0.283 31.72 0.917 0.180
3DGS [16] 24.64 0.731 0.234 30.41 0.920 0.189
SuGaR [11] 22.93 0.629 0.356 29.43 0.906 0.225
Mip-Splatting [47] 24.65 0.729 0.245 30.90 0.921 0.194
2DGS [12] 24.21 0.709 0.276 30.10 0.913 0.211
GS-Pull [55] 23.76 0.703 0.278 30.78 0.925 0.182
GOF [48] 24.82 0.750 0.202 30.79 0.924 0.184
Ours 24.78 0.743 0.211 30.83 0.930 0.166
Table 3: Quantitative Results on Mip-NeRF 360 [3] Dataset. Our method achieves comparable NVS results, especially in the indoor scenes in terms of LPIPS.

### v4

@@ 3.3 Optimization
Geometry Cues.
[…]
$\mathcal{L}_{D}=\sum|({a/(s\times\hat{Z}+t)+b-\bar{D}}|,\ \ \mathcal{L}_{N}=\sum(1-\mathbf{\hat{N}}\cdot\mathbf{\bar{N}}).$ (6)
where $s,t$ and $a,b$ represent learnable scale and shift parameters, and $\mathbf{\bar{N}}$ is rendered Gaussian normal. Therefore, the overall geometry regularization is $\mathcal{L}_{geo}=\lambda_{D}\mathcal{L}_{D}+\lambda_{N}\mathcal{L}_{N}$.
@@
Multi-View Consistent 3D Regularization.
To train this hybrid model, enforcing all Gaussians to satisfy $\bar{s}_{i}=0$ (leading to $\alpha_{i}=1$) degrades rendering quality, as high-quality renderings require opacity variation, and sparse SDF supervision at Gaussian positions is insufficient. Dense supervision with monocular cues like MonoSDF [Yu et al.(2022)Yu, Peng, Niemeyer, Sattler, and Geiger] suffers from scale ambiguity and lacks multi-view consistency, while methods like 3DGSR [Lyu et al.(2024)Lyu, Sun, Huang, Wu, Yang, Chen, Pang, and Qi] and GSDF [Yu et al.(2024a)Yu, Lu, Xu, Jiang, Xiangli, and Dai] only offer 2D supervision through rendered Gaussian depth maps. Our approach leverages monocular cues and the 3D Gaussians’ advantages, where the Gaussian-rendered depth $\bar{D}$ ensures multi-view consistency and provides reliable 3D SDF supervision via back-projection.
More specifically, we first sample $M$ pixels $\{\mathbf{u}=\mathbf{u}_{i}\ |\ \mathbf{u}_{i}\in\mathbb{R}^{2},{\text{for}\ i=1,2,...,M\}}$ from the rendered Gaussian depth map $\bar{D}$. For each ray casts from a pixel, we sample $K=K_{n}+K_{f}$ points $\{t_{k}\}$ from the camera center to the surface, where $K_{n}$ means near surface samples and $K_{f}$ means free space samples. SDF values of each point can be queried as $\{\bar{s}_{k}\}$. For near surface samples within the truncation region $tr$, i.e. $S_{tr}=\{|\bar{D}(\mathbf{u})-t_{k}|\leq tr\}$, we use the distance between the sampled point $t_{k}$ and its surface $\bar{D}(\mathbf{u}_{j})$ as an approximation of ground truth SDF value for supervision, and for points that are far from the surface $S_{fs}=\{\bar{D}(\mathbf{u})-t_{k}>tr\}$, we apply a free-space supervision:
@@
Multi-Resolution Regularization.
Unlike vanilla 3DGS [Kerbl et al.(2023)Kerbl, Kopanas, Leimkühler, and Drettakis] with L1/SSIM losses and opacity-based pruning, our method links opacity to spatial position, eliminating Opacity Resets but risking Gaussian redundancy. Instead of frequency-based strategies [Zhang et al.(2024b)Zhang, Zhan, Xu, Lu, and Xing] that may blur structure, we introduce wavelet-based multi-resolution supervision: $\mathbf{C}_{l}=\mathcal{W}(\mathbf{C},l)$, progressively refining detail while preserving structure. $\mathbf{C}$ is ground truth color images, $\mathcal{W}$ represents wavelet transform and $l$ denotes the level of transformation. This stabilizes training and enables coarse-to-fine Gaussian distribution, preventing redundant Gaussians. The photometric loss is $\mathcal{L}_{p}=0.8|\mathbf{C}_{l}-\bar{\mathbf{C}}|+0.2\cdot SSIM(\mathbf{C}_{l},\bar{\mathbf{C}})$.
Objective Function.
During training, we jointly optimize explicit Gaussians and implicit SDF. We apply distortion $\mathcal{L}_{d}$ and depth-normal $\mathcal{L}_{n}$ regularization as defined in GOF [Yu et al.(2024c)Yu, Sattler, and Geiger]. The overall objective function is defined as: $\mathcal{L}=\mathcal{L}_{p}+\mathcal{L}_{geo}+\mathcal{L}_{sdf}+\lambda_{d}\mathcal{L}_{d}+\lambda_{n}\mathcal{L}_{n}$.
Pruning.
In our framework, opacity is tied to spatial position via SDF. We adopt a geometry-aware pruning strategy that removes Gaussians with $\bar{s}_{i}>tr$, replacing conventional opacity-based pruning with a surface-aware, physically grounded criterion.
24 37 40 55 63 65 69 83 97 105 106 110 114 118 122 Mean
NeRF [Mildenhall et al.(2021)Mildenhall, Srinivasan, Tancik, Barron, Ramamoorthi, and Ng] 1.90 1.60 1.85 0.58 2.28 1.27 1.47 1.67 2.05 1.07 0.88 2.53 1.06 1.15 0.96 1.49
VolSDF [Yariv et al.(2021)Yariv, Gu, Kasten, and Lipman] 1.14 1.26 0.81 0.49 1.25 0.70 0.72 1.29 1.18 0.70 0.66 1.08 0.42 0.61 0.55 0.86
NeuS [Wang et al.(2021)Wang, Liu, Liu, Theobalt, Komura, and Wang] 1.00 1.37 0.93 0.43 1.10 0.65 0.57 1.48 1.09 0.83 0.52 1.20 0.35 0.49 0.54 0.84
N-angelo [Li et al.(2023)Li, Müller, Evans, Taylor, Unberath, Liu, and Lin] 0.49 1.05 0.95 0.38 1.22 1.10 2.16 1.68 1.78 0.93 0.44 1.46 0.41 1.13 0.97 1.07
SuGaR [Guédon and Lepetit(2024)] 1.47 1.33 1.13 0.61 2.25 1.71 1.15 1.63 1.62 1.07 0.79 2.45 0.98 0.88 0.79 1.33
2DGS [Huang et al.(2024)Huang, Yu, Chen, Geiger, and Gao] 0.48 0.91 0.39 0.39 1.01 0.83 0.81 1.36 1.27 0.76 0.70 1.40 0.40 0.76 0.52 0.80
GSurfel [Dai et al.(2024)Dai, Xu, Xie, Liu, Wang, and Xu] 0.66 0.93 0.52 0.41 1.06 1.14 0.85 1.29 1.53 0.79 0.82 1.58 0.45 0.66 0.53 0.88
3DGSR [Lyu et al.(2024)Lyu, Sun, Huang, Wu, Yang, Chen, Pang, and Qi] 0.68 0.84 0.70 0.39 1.16 0.87 0.77 1.48 1.26 0.87 0.69 0.80 0.42 0.64 0.60 0.81
GS-Pull [Zhang et al.(2025)Zhang, Liu, and Han] 0.51 0.56 0.46 0.39 0.82 0.67 0.85 1.37 1.25 0.73 0.54 1.39 0.35 0.88 0.42 0.75
GOF [Yu et al.(2024c)Yu, Sattler, and Geiger] 0.50 0.82 0.37 0.37 1.12 0.74 0.73 1.18 1.29 0.68 0.77 0.90 0.42 0.66 0.49 0.74
Table 1: Quantitative Comparison on the DTU [Jensen et al.(2014)Jensen, Dahl, Vogiatzis, Tola, and Aanæs]. We show the Chamfer distance. For N-angelo [Li et al.(2023)Li, Müller, Evans, Taylor, Unberath, Liu, and Lin], we report the results from UniSDF [Wang et al.(2023)Wang, Rakotosaona, Niemeyer, Szeliski, Pollefeys, and Tombari] reproduction and we show the vanilla results in supp. mat..
Figure 3: Surface Reconstruction on the DTU [Jensen et al.(2014)Jensen, Dahl, Vogiatzis, Tola, and Aanæs]. We show normal maps from the reconstructed meshes.
@@ 4.1 Experimental Settings
Datasets.
We comprehensively evaluate our method on three public datasets: DTU dataset [Jensen et al.(2014)Jensen, Dahl, Vogiatzis, Tola, and Aanæs] which consists of indoor object scans; Tanks and Temples [Knapitsch et al.(2017)Knapitsch, Park, Zhou, and Koltun] which features six unbounded outdoor scenes; and Mip-NeRF 360 [Barron et al.(2022)Barron, Mildenhall, Verbin, Srinivasan, and Hedman] for measuring novel view synthesis quality.
Baselines.
Metrics.
We follow common practice and report surface accuracy as Chamfer Distance and F1-score on DTU and Tanks and Temples, respectively. We measure the visual fidelity of the synthesized novel views with PSNR, SSIM and LPIPS [Zhang et al.(2018)Zhang, Isola, Efros, Shechtman, and Wang] on Mip-NeRF 360.
Implementation.
We perform single GPU training (NVIDIA 3090) and use by default $\lambda_{d}=100,\lambda_{n}=0.05,\lambda_{ns}=1000,\lambda_{fs}=10,\lambda_{depth}=0.05,\lambda_{normal}=0.1$ and $\beta=100$. We train 30000 iterations like other methods and do not require extra training for SDF. For more implementation details, please refer to supp. mat..
@@
NeuS [Wang et al.(2021)Wang, Liu, Liu, Theobalt, Komura, and Wang] GeoNeus [Fu et al.(2022)Fu, Xu, Ong, and Tao] N-angelo [Li et al.(2023)Li, Müller, Evans, Taylor, Unberath, Liu, and Lin] SuGaR [Guédon and Lepetit(2024)] 2DGS [Huang et al.(2024)Huang, Yu, Chen, Geiger, and Gao] GOF [Yu et al.(2024c)Yu, Sattler, and Geiger] Ours
@@
Mean 0.38 0.35 0.50 0.19 0.30 0.46 0.47
Table 2: Quantitative Results on the Tanks and Temples [Knapitsch et al.(2017)Knapitsch, Park, Zhou, and Koltun].We show the F1-score.
4.2 Geometry Evaluation
We evaluate our method on the DTU dataset [Jensen et al.(2014)Jensen, Dahl, Vogiatzis, Tola, and Aanæs], achieving the lowest Chamfer Distance and outperforming all 3DGS-based and implicit methods in quality (Table 1). For N-angelo [Li et al.(2023)Li, Müller, Evans, Taylor, Unberath, Liu, and Lin], we report results from [Wang et al.(2023)Wang, Rakotosaona, Niemeyer, Szeliski, Pollefeys, and Tombari] due to unverified original values. As shown in Figure 3, our method yields smoother, more complete surfaces—especially in sparse and flat regions—compared to noisy outputs from 2DGS [Huang et al.(2024)Huang, Yu, Chen, Geiger, and Gao] and GOF [Yu et al.(2024c)Yu, Sattler, and Geiger]. On the Tanks and Temples dataset [Knapitsch et al.(2017)Knapitsch, Park, Zhou, and Koltun], as shown in Table 2, our method matches the performance of leading implicit approaches [Li et al.(2023)Li, Müller, Evans, Taylor, Unberath, Liu, and Lin] while cutting training time from 12+ to around 3 hours. We also slightly outperform GOF [Yu et al.(2024c)Yu, Sattler, and Geiger], with faster mesh extraction and similar training speed (Table 4). Figure 4 shows our superior handling of flat and transparent surfaces. Additional results on Mip-NeRF 360 [Barron et al.(2022)Barron, Mildenhall, Verbin, Srinivasan, and Hedman], DTU [Jensen et al.(2014)Jensen, Dahl, Vogiatzis, Tola, and Aanæs], and Tanks and Temples [Knapitsch et al.(2017)Knapitsch, Park, Zhou, and Koltun] are in the supp. mat..
Figure 4: Surface Reconstruction on the Tanks and Temples [Knapitsch et al.(2017)Knapitsch, Park, Zhou, and Koltun]. We show rendered normal maps from reconstructed meshes.
Metric NeRF [Mildenhall et al.(2021)Mildenhall, Srinivasan, Tancik, Barron, Ramamoorthi, and Ng] InstantNGP [Müller et al.(2022)Müller, Evans, Schied, and Keller] MipNeRF360 [Barron et al.(2022)Barron, Mildenhall, Verbin, Srinivasan, and Hedman] 3DGS [Kerbl et al.(2023)Kerbl, Kopanas, Leimkühler, and Drettakis] SuGaR [Guédon and Lepetit(2024)] Mip-Splat [Yu et al.(2024b)Yu, Chen, Huang, Sattler, and Geiger] 2DGS [Huang et al.(2024)Huang, Yu, Chen, Geiger, and Gao] GS-Pull [Zhang et al.(2025)Zhang, Liu, and Han] GOF [Yu et al.(2024c)Yu, Sattler, and Geiger] Ours
PSNR $\uparrow$ 24.15 26.03 28.10 27.53 26.18 27.78 27.16 27.27 27.81 27.81
SSIM $\uparrow$ 0.624 0.723 0.804 0.826 0.768 0.825 0.811 0.814 0.837 0.836
LPIPS $\downarrow$ 0.443 0.294 0.232 0.212 0.291 0.220 0.244 0.230 0.193 0.189
Table 3: Quantitative Results on Mip-NeRF 360 [Barron et al.(2022)Barron, Mildenhall, Verbin, Srinivasan, and Hedman]. We evaluate the quality of rendered images against ground-truth images using PSNR, SSIM, and LPIPS.